3+sqrt(z-7)=sqrt(z+8)
Answer by nyc_function(2741)

3+sqrt(z-7)=sqrt(z+8)
Square both sides.
(3+sqrt(z-7)^2 = (sqrt(z+8)^2

9 + z - 7 = z + 8
z + 2 = z + 8
At this point, we need to combine like terms. However, when this is done, we can clearly see that z will cancel out.
LOOK:
z - z = 8 - 2
0 = 6
This does not make sense since zero does not = 6.

What is the answer?

The answer, and this happens sometimes with radicals, is NO SOLUTION.

So, there is no solution to this radical equation.